Raina is saving money to buy a bike that $125 costs . She has $60 and will save an additional $5 each week. In how many weeks will she have enough money to buy the bike?
Answer by math_tutor2020(3816)


w = number of weeks
5w = additional amount saved at $5 per week
5w+60 = total amount after adding the initial $60
5w+60 = 125
5w+60-60 = 125-60
5w = 65
5w/5 = 65/5
w = 13 weeks is the final answer